solve (x^3+3xy^2)p+(y^3+3x^2y)q=(x+y)^2z


1
Expert's answer
2021-05-23T23:18:02-0400

Given equation is-

"(x^3+3xy^2)p+(y^3+3x^2y)q=(x+y)^2z"


"\dfrac{dx}{x^3+3xy^2}=\dfrac{dy}{y^3+3x^2y}=\dfrac{dz}{z(x+y)^2}"


"\dfrac{dy}{dx}=\dfrac{y^3+3x^2y}{x^3+3xy^2}"


This is a homogeneous differential equation.


Let "\dfrac{y}{x}=v, y=xv, \dfrac{dy}{dx}=v+x(\dfrac{dv}{dx})"


Then:

"v+x\dfrac{dv}{dx}=\dfrac{(xv)^3+3x^3v}{x^3+3x^3v^2}=\dfrac{v^3+3v}{1+3v^2}"


"x\dfrac{dv}{dx}=\dfrac{v^3+3v}{1+3v^2}-v=\dfrac{4v^3+2v}{1+3v^2}"


"\dfrac{1}{2}\intop\dfrac{1+3v^2}{v(2v^2+1)}dv=\intop \dfrac{dx}{x}"


"\dfrac{1+3v^2}{v(2v^2+1)}=\dfrac{A}{v}+\dfrac{Bv+C}{2v^2+1}"


"A(2v^2+1)+v(Bv+C)=1+3v^2"

"2A+B=3"

"A=1"

"C=0"

"B=3-2A=1"


"\dfrac{1-3v^2}{v(2v^2+1)}=\dfrac{1}{v}+\dfrac{v}{2v^2+1}"


"\intop\dfrac{v}{2v^2+1}dv=\dfrac{1}{2}\intop\dfrac{d(v^2)}{2v^2+1}dv=\dfrac{ln(2v^2+1)}{4}"


"\dfrac{lnv}{2}-\dfrac{5ln(2v^2+1)}{8}=lnx+lnc_1"


"ln(\dfrac{v^{1/2}}{(2v^2+1)^{5/8}})=ln(c_1x)"


"\dfrac{(y/x)^{1/2}}{x(2(y/x)^2+1)^{5/8}}=c_1"



"\dfrac{dx/x+dy/y+dz/z}{6x^2}=-\dfrac{3(dx/x-dy/y+dz/z)}{6y^2}=\dfrac{dz}{z(x+y)^2}"


"\dfrac{dx/x+dy/y+dz/z-3(dx/x-dy/y+dz/z)}{6x^2+6y^2}=\dfrac{3dz/z}{6(x+y)^2}"



"\dfrac{dx}{x}+\dfrac{dy}{y}+\dfrac{dz}{z}-3(\dfrac{dx}{x}-\dfrac{dy}{y}+\dfrac{dz}{z})=3\dfrac{dz}{z}"


"4\dfrac{dy}{y}-2\dfrac{dx}{x}=5\dfrac{dz}{z}"


"4lny-2lnx=5lnz+lnc_2"


"ln(\dfrac{y^4}{x^2})=ln(c_2z^5)"


"\dfrac{y^4}{x^2z^5}=c_2"


The general integral:


"\phi(\dfrac{(y/x)^{1/2}}{x(2(y/x)^2+1)^{5/8}}, \dfrac{y^4}{x^2z^5})=0"